# Thermal Noise Measurement Below the Standard Quantum Limit

### v1

@@ I.1 Thermal Noise
$S_{\mathrm{th}}(\omega)=\sum_{k=0}^{n}\frac{4k_{B}T\omega_{k}\phi}{m_{k}[(\omega^{2}_{k}-\omega^{2})^{2}+\omega^{2}\omega^{2}_{k}\phi^{2}]}.$ (1)
@@ III Results and discussion
The uncertainties of the measured and modeled spectrum were accounted for as follows. The primary uncertainty of $\sim 1\%$ identified in the $S_{0}$ spectrum results from the uncertainty in the calibration from $\mathrm{V^{2}/Hz}$ into $\mathrm{m^{2}/Hz}$, just as in [1]. Uncertainty in the SN and uncorrelated electronics noise, given by the average fluctuations between measurements of $S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}\approx S_{LL}-S_{LM}$ (see Appendix A), results in additional uncertainty in the measurement of $S_{1}$. The uncertainties of modal mass and optical coupling to the cavity also result in uncertainty in our modeled noise spectrum. The contributing parameters are the mass of the oscillator, the temperature recorded while the cryopump is turned off, and the input power of the light pumping the OM cavity.
@@ III.1 Thermal Noise Measurement
Fig. 2 displays measurements in two columns: the first presents the amplitude spectral density (ASD) in units of $\mathrm{m/\sqrt{Hz}}$ before the SN is removed, normalized to the free mass SQL ($\sqrt{S_{0}/S_{\mathrm{SQL}}}$, where $S_{0}$ is given by Eq. 6 and $S_{\mathrm{SQL}}$ is $S_{\mathrm{SQL}}=\frac{2\hbar}{m\omega^{2}}$). The second column presents the ASD after SN and the uncorrelated electronics noise is removed, also normalized to the SQL ($\sqrt{S_{1}/S_{\mathrm{SQL}}}$, with $S_{1}$ given by Eq. 7). All datasets display unmodeled lines throughout the measurements, which is most likely the result of electronics noise. There is also some separation between the modeled and measured noise, which is more pronounced before the uncorrelated noise is removed, suggesting that not all optical losses present in the cavity are accounted for, or that uncorrelated electronics noise is not negligible. All measurements show clear evidence that SN is removed in the second column. The 50 kHz OS frequency spectrum falls 2.2 dB below the SQL before the uncorrelated noise is removed. Much of the frequency range considered here exceeds the 50 kHz OS frequency present in this measurement, meaning that the cross correlation measurement is unreliable at these frequencies (see discussion in Sec. II.1 and Appendix A). However, we do see clear evidence of the $S_{1}$ spectrum falling $\sim 2$ dB below the SQL at frequencies below the OS frequency.
@@ Appendix A Calibration
In order to remove the LFN and complete the calibration into $\mathrm{m/\sqrt{Hz}}$, a cross power spectral density (CPSD) between $PD_{\mathrm{F}}$ and one of the PDs monitoring the cavity signal was recorded along with the PSD of $PD_{\mathrm{L}}$ and $PD_{\mathrm{F}}$. While either $PD_{\mathrm{L}}$ or $PD_{\mathrm{M}}$ could be used, $PD_{\mathrm{L}}$ was used here. The derivation of a conversion factor, used to calibrate the spectra into the appropriate units is documented in [1, 30]. In order to remove the SN, a second CPSD and the corresponding PSDs are measured between $PD_{\mathrm{L}}$ and $PD_{\mathrm{M}}$. To give a full account of the calculations performed on this data, we first consider the power fluctuations measured on $PD_{\mathrm{F}}$, given by:
$\hat{P}_{\mathrm{F}}=A(\hat{v}_{\mathrm{v,f}}+\hat{v}_{\mathrm{f}})$ (10)
@@
$=\frac{1}{1+G_{\mathrm{fb}}}\left[\frac{1}{\sqrt{2}}B(\hat{v}_{\mathrm{v,0}}+\hat{v}_{\mathrm{v,bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{L}}\right],$
Here $\hat{P}_{\mathrm{L,0}}$ represents the power fluctuations measured on the respective PDs before the feedback is applied, as described in Eq. 2, $G_{\mathrm{fb}}$ is the open loop gain of the feedback, and $B$ is the field amplitude of the light leaving the cavity.
@@
$S_{\mathrm{LL}}=\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{0}}+S_{\mathrm{bs}}+S_{\mathrm{cl}}+S_{\mathrm{el,L}}\right)$
$=\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{rp}}+S_{\mathrm{sn,L}}+S_{\mathrm{f}}+S_{\mathrm{th}}+S_{\mathrm{el,L}}\right).$ (13)
Here $S_{\mathrm{f,fm}}$ is the LFN measured after the MZ interferometer, before the optical spring gain affects the noise source; $S_{\mathrm{sn,F}}$ is the SN and $S_{\mathrm{el,F}}$ is the electronics noise measured on $PD_{\mathrm{F}}$; $S_{\mathrm{sn,L}}$ and $S_{\mathrm{el,L}}$ are the respective SN and electronics noise measured on $PD_{\mathrm{L}}$ before the feedback is applied to the photodetector; last, $S_{\mathrm{rp}}$ is the noise resulting from the QRPN and $S_{\mathrm{cl}}=S_{\mathrm{f}}+S_{\mathrm{th}}$ is the classical noise, with contributions from the LFN ($S_{\mathrm{f}}$) and TN ($S_{\mathrm{th}}$). These spectra represent the noise after it is amplified by the optical spring gain $G_{\mathrm{os}}$. Expressing the power spectrum explicitly in terms of $G_{\mathrm{os}}$,
$S_{\mathrm{LL}}=\left(\frac{G_{\mathrm{os}}}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{rp,fm}}+S_{\mathrm{f,fm}}+S_{\mathrm{th,fm}}\right)$ (14)
$+\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}\right).$
@@
$S_{0}$ $=K^{2}_{\mathrm{m}}S_{\mathrm{FF}}\left(\frac{1-C}{C}\right)$ (18)
@@
$\hat{P}_{\mathrm{M}}$ $=\hat{P}_{\mathrm{M,0}}-G_{\mathrm{fb}}\hat{P}_{\mathrm{L}}$ (19)
$=\frac{1}{\sqrt{2}}B(\hat{v}_{\mathrm{v,0}}-\hat{v}_{\mathrm{v,bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{M}}$
$-\frac{G_{\mathrm{fb}}}{1+G_{\mathrm{fb}}}\left[\frac{1}{\sqrt{2}}iB^{*}(\hat{v}_{\mathrm{v,0}}+\hat{v}_{\mathrm{v,bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{L}}\right].$
@@
$S_{\mathrm{exs}}=2\frac{G_{\mathrm{fb}}}{(1+G_{\mathrm{fb}})^{2}}S_{\mathrm{b}}+\frac{G_{\mathrm{fb}}}{(1+G_{\mathrm{fb}})^{2}}S_{\mathrm{el,L}}$ (23)
and $S_{0}-S_{\mathrm{b}}=G_{\mathrm{os}}S_{\mathrm{rp,fm}}$.
The QRPN and TN in the measurement can be evaluated by substituting $S_{\mathrm{LM}}$ in place of $S_{\mathrm{LL}}$ in equation Eq. 16 then follow the steps ending in Eq. 18. This results in the equation calibrated into $m^{2}/Hz$:
$S_{1}=S_{rp,fm}+S_{th,fm}+\frac{G_{\mathrm{fb}}}{G^{2}_{\mathrm{os}}}(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}).$ (24)

### v2

@@ I.1 Thermal Noise
$S_{\mathrm{th}}(\omega)=\sum_{k=0}^{n}\frac{4k_{B}T\omega_{k}^{2}\phi}{m_{k}\omega[(\omega^{2}_{k}-\omega^{2})^{2}+\omega^{4}_{k}\phi^{2}]}.$ (1)
@@ III Results and discussion
The uncertainties of the measured and modeled spectrum were accounted for as follows. The primary source of uncertainty of $\sim 1\%$ identified in the $S_{0}$ spectrum results from the uncertainty in the calibration from $\mathrm{V^{2}/Hz}$ into $\mathrm{m^{2}/Hz}$, just as in [1]. Additional variation in $S_{1}$ spectra resulted from variations in the measured SN and uncorrelated electronics noise. A $10\%$ uncertainty in the mass of the mirror, an $8\%$ uncertainty in the temperature, and a $9\%$ uncertainty in the power of the light pumping the cavity all result in uncertainty in our modeled noise spectrum.
@@ III.1 Thermal Noise Measurement
Fig. 2 displays measurements in two columns: the first presents the amplitude spectral density (ASD) in units of $\mathrm{m/\sqrt{Hz}}$ before the SN is removed, normalized to the free mass SQL ($\sqrt{S_{0}/S_{\mathrm{SQL}}}$, where $S_{0}$ is given by Eq. 6, and $S_{\mathrm{SQL}}=\frac{2\hbar}{m\omega^{2}}$). The second column presents the ASD after SN and the uncorrelated electronics noise is removed, also normalized to the SQL ($\sqrt{S_{1}/S_{\mathrm{SQL}}}$, with $S_{1}$ given by Eq. 7). All datasets display unmodeled lines throughout the measurements, which is most likely the result of electronics noise. There is also some separation between the modeled and measured noise, which is more pronounced before the uncorrelated noise is removed, suggesting that not all optical losses present in the cavity are accounted for, or that uncorrelated electronics noise is not negligible. All measurements show clear evidence that SN is removed in the second column. The 50 kHz OS frequency spectrum falls 2.2 dB below the SQL before the uncorrelated noise is removed. Much of the frequency range considered here exceeds the 50 kHz OS frequency present in this measurement, meaning that the cross correlation measurement is unreliable at these frequencies (see discussion in Sec. II.1 and Appendix A). However, we do see clear evidence of the $S_{1}$ spectrum falling $\sim 2$ dB below the SQL at frequencies below the OS frequency.
@@ Appendix A Calibration
In order to remove the LFN and complete the calibration into $\mathrm{m/\sqrt{Hz}}$, a cross power spectral density (CPSD) between $PD_{\mathrm{F}}$ and one of the PDs monitoring the cavity signal was recorded along with the PSD of both $PD_{\mathrm{L}}$ and $PD_{\mathrm{F}}$. While either $PD_{\mathrm{L}}$ or $PD_{\mathrm{M}}$ could be used, $PD_{\mathrm{L}}$ was used here. The derivation of a conversion factor, used to calibrate the spectra into the appropriate units is documented in [1, 30]. In order to remove the SN, a second CPSD and the corresponding PSDs are measured between $PD_{\mathrm{L}}$ and $PD_{\mathrm{M}}$. To give a full account of the calculations performed on this data, we first consider the power fluctuations measured on $PD_{\mathrm{F}}$, given by:
$\hat{P}_{\mathrm{F}}=A(\hat{v}_{\mathrm{v,f}}+\hat{v}_{\mathrm{f}})+\hat{P}^{\mathrm{el}}_{\mathrm{F}}$ (10)
@@
$=\frac{1}{1+G_{\mathrm{fb}}}\left[\frac{1}{\sqrt{2}}B(\hat{v}_{\mathrm{0}}+\hat{v}_{\mathrm{bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{L}}\right].$
Here $\hat{P}_{\mathrm{L,0}}$ represents the power fluctuations measured on the respective PDs before the feedback is applied, $G_{\mathrm{fb}}$ is the open loop gain of the feedback. In the second line B is the field amplitude; $\hat{v}_{\mathrm{0}}$ and $\hat{v}_{\mathrm{bs}}$ are vacuum fluctuations exiting the cavity and coupling into the signal through the unused port of the beamsplitter respectively; $\hat{v}_{\mathrm{cl}}$ results from any classical noise which most notably includes TN, and $\hat{P}^{\mathrm{el}}_{\mathrm{L/M}}$ represents the power fluctuations from uncorrelated electronics noise (e.g. the dark current from the two PDs) just as defined in Eq. 2.
@@
$S_{\mathrm{LL}}$ $=\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{0}}+S_{\mathrm{bs}}+S_{\mathrm{cl}}+S_{\mathrm{el,L}}\right)$ (13)
$=\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{rp}}+S_{\mathrm{sn,L}}+S_{\mathrm{f}}+S_{\mathrm{th}}+S_{\mathrm{el,L}}\right).$
Each element of the Eqs. 12 and 13 comes from a corresponding element of the Eqs. 10 and 11. $S_{\mathrm{f,fm}}$ is the LFN measured after the MZ interferometer, before the optical spring gain affects the noise source; $S_{\mathrm{sn,F}}$ is the SN and $S_{\mathrm{el,F}}$ is the electronics noise measured on $PD_{\mathrm{F}}$; $S_{\mathrm{sn,L}}$ and $S_{\mathrm{el,L}}$ are the respective SN and electronics noise measured on $PD_{\mathrm{L}}$ before the feedback is applied to the photodetector; last, $S_{\mathrm{rp}}$ is the noise resulting from the QRPN and $S_{\mathrm{cl}}=S_{\mathrm{f}}+S_{\mathrm{th}}$ is the classical noise, with contributions from the LFN ($S_{\mathrm{f}}$) and TN ($S_{\mathrm{th}}$). These spectra represent the noise after it is amplified by the optical spring gain $G_{\mathrm{os}}$. Expressing the power spectrum explicitly in terms of $G_{\mathrm{os}}$,
$S_{\mathrm{LL}}$ $=\left(\frac{G_{\mathrm{os}}}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{rp,fm}}+S_{\mathrm{f,fm}}+S_{\mathrm{th,fm}}\right)$
$+\left(\frac{1}{1+G_{\mathrm{fb}}}\right)^{2}\left(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}\right).$ (14)
@@
$S_{0}$ $=K^{2}_{\mathrm{m}}S_{\mathrm{FF}}\left(\frac{1-C}{C}\right)$
@@
$\hat{P}_{\mathrm{M}}$ $=\hat{P}_{\mathrm{M,0}}-G_{\mathrm{fb}}\hat{P}_{\mathrm{L}}$
$=\frac{1}{\sqrt{2}}B(\hat{v}_{\mathrm{0}}-\hat{v}_{\mathrm{bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{M}}$
$-\frac{G_{\mathrm{fb}}}{1+G_{\mathrm{fb}}}\left[\frac{1}{\sqrt{2}}iB^{*}(\hat{v}_{\mathrm{0}}+\hat{v}_{\mathrm{bs}}+\hat{v}_{\mathrm{cl}})+\hat{P}^{\mathrm{el}}_{\mathrm{L}}\right].$
@@
$S_{\mathrm{exs}}=\frac{G_{\mathrm{fb}}}{(1+G_{\mathrm{fb}})^{2}}S_{\mathrm{sn,L}}+\frac{G_{\mathrm{fb}}}{(1+G_{\mathrm{fb}})^{2}}S_{\mathrm{el,L}},$ (23)
$S_{0}-S_{\mathrm{b}}=G_{\mathrm{os}}S_{\mathrm{rp,fm}},$ and $2S_{\mathrm{b}}=S_{\mathrm{sn,L}}$.
We can then arrive at an upper bound of the shotnoise present in $S_{0}$:
$\left(S_{\mathrm{LL}}-S_{\mathrm{LM}}\right)\times\left(\frac{1+G_{\mathrm{fb}}}{G_{\mathrm{os}}}\right)^{2}$ $=\left(\frac{1}{G_{\mathrm{op}}}\right)^{2}\left(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}\right)$ (24)
$+\left(\frac{1+G_{\mathrm{fb}}}{G_{\mathrm{os}}}\right)^{2}\times S_{\mathrm{exs}}.$
Note, to evaluate equation Eq. 24, we use
$\left(\frac{G_{\mathrm{os}}}{1+G_{\mathrm{fb}}}\right)=\frac{S_{\mathrm{FL}}}{S_{\mathrm{FF}}}.$ (25)
Finally subtracting Eq. 24 from Eq. A we find the spectra, calibrated into $m^{2}/Hz$:
$S_{1}=S_{\mathrm{rp,fm}}+S_{\mathrm{th,fm}}-\frac{G_{\mathrm{fb}}}{G^{2}_{\mathrm{os}}}(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}}).$ (26)
Here we see that our meaurement is not exact by the quantity $\frac{G_{\mathrm{fb}}}{G^{2}_{\mathrm{os}}}(S_{\mathrm{sn,L}}+S_{\mathrm{el,L}})$, adding some uncertainty to the measurement. However this quantity is negligible close to the optical spring frequency: $G_{\mathrm{fb}}/G^{2}_{\mathrm{os}}<<0$.